### v1

# Towards deployment-centric multimodal AI beyond vision and language

Abstract
Multimodal artificial intelligence (AI) integrates diverse types of data via machine learning to improve understanding, prediction, and decision-making across disciplines such as healthcare, science, and engineering. However, most multimodal AI advances focus on models for vision and language data, while their deployability remains a key challenge. We advocate a deployment-centric workflow that incorporates deployment constraints early to reduce the likelihood of undeployable solutions, complementing data-centric and model-centric approaches. We also emphasise deeper integration across multiple levels of multimodality and multidisciplinary collaboration to significantly broaden the research scope beyond vision and language. To facilitate this approach, we identify common multimodal-AI-specific challenges shared across disciplines and examine three real-world use cases: pandemic response, self-driving car design, and climate change adaptation, drawing expertise from healthcare, social science, engineering, science, sustainability, and finance. By fostering multidisciplinary dialogue and open research practices, our community can accelerate deployment-centric development for broad societal impact.
Introduction
Data drives discovery in the 21st century wang2023scientific. From satellites monitoring climate change to social media capturing human behaviour, the variety and volume of information have never been greater. Each type of data, or modality, offers unique insights, but unimodal approaches often fall short of achieving robust or generalisable performance. Autonomous vehicles relying solely on visual data struggle with object detection in low-light or adverse weather conditions kabir2025terrain. During the COVID-19 pandemic, relying solely on RT-PCR data for diagnosing SARS-CoV-2 infection led to frequent false negatives, hindering timely interventions woloshin2020false. Instead, integrating information from multiple modalities can reveal patterns and solutions that unimodal approaches miss Zhao2024DeepMD.
Multimodal artificial intelligence (AI) leverages multimodal data for better understanding complex systems through machine learning baltruvsaitis2018multimodal; ektefaie2023multimodal; xu2023multimodal; liang2024foundations. Its promise is evident in multidisciplinary applications such as pandemic response. For example, in healthcare, combining medical imaging, genomic sequencing, and epidemiological data can enhance diagnoses, inform treatment strategies, and support disease prevention acosta2022multimodal; kline2022multimodal; krones2025review. In science, fusing genetic and protein structure data holds promise for advancing vaccine development notin2024machine. In engineering, integrating textual specifications with spatial data can improve product design and manufacturing song2024multi, which could extend to optimising ventilator production. Across other disciplines, such as sustainability, finance, and social science, multimodal AI offers the potential to provide deeper insights and actionable strategies ofodile2024predictive; quatrini2021challenges; gupta2021emotion; anshul2023multimodal. However, many current approaches face challenges in deployment due to data limitations, integration complexities, and practical constraints roberts2021common.
While multimodal AI receives increasing attention and holds great promise, research has primarily focused on vision-language models bordes2024introduction; zhang2024vision, leaving other modalities, such as tabular and time-series data, and related disciplines underexplored vanposition (Fig. 1 and Supplementary Fig. S2). Real-world challenges, such as pandemic response, call for AI capable of integrating diverse types of data through multidisciplinary collaboration, bridging the gap between research and application.
This Perspective outlines a deployment-centric framework for multimodal AI that incorporates deployment constraints early and addresses challenges specific to multimodal integration across disciplines. We first present a general workflow for developing deployment-ready multimodal AI systems. We then examine three data-intensive, cross-disciplinary use cases, pandemic response, self-driving car design, and climate change adaptation, to illustrate common barriers and actionable strategies. By exploring underrepresented modalities and fostering open, interdisciplinary collaboration, we aim to broaden the impact of multimodal AI beyond vision and language.
Figure 1: Trends in multimodal AI research and dominance of vision and language. a, Yearly growth of multimodal AI preprints on arXiv by modality, showing a steady increase over time and the dominance of vision and language. b, Breakdown of modality pairs in multimodal AI preprints on arXiv in 2024, revealing that over half the studies involve vision and language, followed by vision and others (19.9%), language and others (18.0%), and other modality combinations (11.1%). This analysis highlights the most common pairwise modality combinations and shows that those involving vision or language dominate, reaching 88.9%. For clarity and space efficiency, only modality pairs exceeding 1.5% are annotated. See Supplementary Fig. S2 for a more detailed illustration and analysis of the multimodal AI landscape, including statistics on triple and quadruple modality combinations.
Deployment-centric multimodal AI system development
Traditionally, multimodal AI research has been model-centric baltruvsaitis2018multimodal; ektefaie2023multimodal; xu2023multimodal, focusing on developing new models to outperform existing ones on standardised benchmarks and datasets. The rise of generative AI jo2023promise models such as ChatGPT brown2020language; achiam2023gpt resulted in a shift towards data-centric approaches seedat2023navigating; liu2024visual; li2024llava, emphasising data resources and quality for better model performance. However, the gap between the high expectations and limited real-world impacts roberts2021common; paleyes2022challenges indicates a pressing need for deployment-centric approaches that prioritise real-world applicability, user needs, and ethical considerations, ensuring that AI innovations are both novel and practical, with a positive impact. We advocate a deployment-centric workflow for advancing multimodal AI from research to scalable solutions, built on ideas from the machine learning operations lifecycle guidelines kreuzberger2023machine and technology readiness levels for machine learning systems lavin2022technology.
Figure 2: Deployment-centric multimodal AI: workflow and challenges. a, Deployment-centric multimodal AI workflow designed to meet real-world needs through a structured three-stage process covering planning, development, and deployment, with iterative assessments and updates. In particular, the planning stage considers deployment constraints early to ensure alignment with real-world settings and practical needs. This workflow incorporates multidisciplinary collaboration at all stages to ensure discipline-specific knowledge informs and enhances AI system development. Moreover, we consider multiple levels of data multimodality, from data types and subtypes to views and fidelities, as illustrated with a healthcare example. This broader definition of multimodality offers new perspectives and rich options for leveraging the benefits of multimodality. The system development stage has five steps similar to a standard machine learning pipeline, where the data-centric and model-centric approaches seedat2023navigating are indicated in the figure to highlight their differences from the deployment-centric approach. b, Five multimodal-AI-specific challenges shared across multiple disciplines and real-world applications.
Figure 2a illustrates our deployment-centric workflow in three stages: planning, development, and deployment. The planning stage focuses on defining the problem, determining the suitability of multimodal AI over unimodal AI, understanding real-world constraints, and formulating AI tasks grounded in realistic assumptions. The development stage builds the multimodal AI system to learn predictive models from multimodal data, and the deployment stage assesses real-world performance, feeding back insights to improve earlier steps. Guidance from multidisciplinary collaboration (the top right of Fig. 2a) informs the entire workflow, ensuring that technical, ethical, and societal perspectives are integrated into the process for a robust and responsible system.
Planning for multimodal AI systems
The planning stage (the three blocks on the top left of Fig. 2a) begins with problem definition, which involves clearly articulating the problem’s scope and objectives, and preliminarily assessing whether incorporating multimodal data can offer meaningful advantages, such as improved predictive performance or deeper insights, over unimodal approaches huang2021makes. Beyond vision (image, video) and language (text), key modalities include audio, numeric time series (e.g. financial data sequences), sensor signals (e.g. wearable physiological measurements), and spatial (geolocation), tabular (structured data such as clinical records), and graph-based (relationships or networks) data. Furthermore, multimodality exists at multiple levels: multiple data types, subtypes, views, and fidelities meng2021multi; Penwarden2022MultifidelityPINNs (the middle right of Fig. 2a). For example, in healthcare, multimodal data can range from different types, e.g. electronic health records (EHRs), medical images, and omics data zitnik2019machine; lunke2023integrated, to subtypes within a data type, e.g. X-ray, magnetic resonance imaging (MRI), and ultrasound within the imaging modality, multiple views of the same data (sub)type, e.g. anteroposterior, lateral, and oblique views of X-ray, and multiple fidelities, e.g. high, low, and adaptive fidelities of anteroposterior X-ray. Under this broad definition, the potential number of modalities can expand significantly as more levels are considered. Understanding these modalities early helps select those suited to the problem’s complexity and capable of offering insights beyond unimodal data.
After establishing the potential benefits of multimodality, we move on to deployment constraints, which involves examining the AI system’s application context, including user needs, data availability, regulatory compliance wu2021medical, ethical considerations, societal impact, and economic trade-offs between high- and low-cost data modalities. This step ensures that the chosen modalities and the resulting multimodal AI system are not only technically feasible but also viable within the intended deployment environments. Understanding these constraints early informs task formulation, helping align the solution with real-world requirements and making it more effective and responsible.
The final step, task formulation, translates the defined problem and constraints into specific AI tasks, specifying the inputs, outputs, and evaluation metrics required to meet the objectives set during problem definition. This provides a clear development roadmap, ensuring the multimodal AI system is well-positioned to meet both technical and practical needs.
Multimodal AI system development
Developing multimodal AI systems (the lower left of Fig. 2a) parallels standard machine learning system development but introduces added complexities due to the integration of diverse data modalities. We have identified five multimodal-AI-specific challenges (Fig. 2b). Modality incompleteness occurs when one or more modalities are missing during training or deployment, necessitating the development of highly flexible or generative models. Multimodal heterogeneity deals with varying formats, scales, and data structures, necessitating careful integration strategies to ensure interoperability. Cross-modality alignment ensures that the timing and/or meaning of data from different sources are correctly aligned for coherency and consistency, whether temporal (e.g. matching timestamps) or semantic (e.g. identifying data representing the same concept across modalities). Modality complementarity ensures that different modalities contribute complementary information to enhance performance, as more modalities may not improve results, e.g. if they add noise or redundancy. Finally, multimodal privacy risk zhao2024survey; pranjal2023toward arises when independently anonymised datasets become identifiable again (re-identification) through their fusion, thereby revealing sensitive information and necessitating robust privacy-preserving techniques.
The development process consists of five key steps: data collection, data curation, multimodal learning, evaluation, and interpretation (the lower right of Fig. 2a). Each step plays a key role in addressing the multimodal-AI-specific challenges described above.
Data collection and curation build on the modality choices identified during planning, gathering relevant data from multiple sources and preparing it for AI model training. Diverse and reliable data sources ensure a high-quality, comprehensive representation of the problem space. Once collected, the data undergoes standard curation, such as wrangling, cleaning, annotation, handling missing values, and quality assurance paleyes2022challenges, tailored to the needs of multimodal AI. For instance, in healthcare, linking multimodal data from EHRs, imaging, and biosignals poses significant challenges due to privacy concerns zhao2024survey and the need for standardisation. The integration and cross-referencing of these modalities can amplify the risk of re-identification, underscoring the importance of addressing multimodal privacy risk through robust privacy-preserving techniques al2019privacy; Wendland:2022vo; che2023MultimodalFL. Moreover, selected modalities should be interoperable and complementary, with their alignment and integration resulting in a cohesive, high-quality dataset suited to both technical and deployment requirements.
Multimodal learning integrates diverse data modalities to leverage their unique strengths for improved predictive performance baltruvsaitis2018multimodal. Traditional fusion strategies, early fusion (combining raw features), intermediate fusion (merging processed features), and late fusion (integrating outputs from independently processed modalities), offer trade-offs between model complexity and the timing of modality interaction. While these methods are well-established, recent advances such as hybrid fusion Zhao2024DeepMD and knowledge distillation Wang2020MultimodalLW provide greater flexibility, enabling the combination of multiple strategies or the transfer of knowledge from complex to simpler models. Techniques such as co-attention mechanisms Yu2017MultimodalFB can further enhance integration and performance by dynamically adapting to the context of cross-modal interactions. The choice of strategy depends on the specific problem and data characteristics, balancing integration benefits with the added complexity of managing multiple modalities.
Evaluation and interpretation support the reliability, effectiveness, and fairness of multimodal AI systems, especially in complex, high-stakes environments. Beyond standard performance metrics such as accuracy, it is important to assess the contribution of individual modalities, the alignment and synergy between them, and the system’s robustness to noise and variability. For example, evaluating safety in multimodal models weidinger2024holistic is key for understanding how biases across modalities may interact or amplify one another luccioni2024stable, potentially undermining system reliability. Interpretation murdoch2019definitions; rudin2019stop ensures that model decisions are understandable and transparent. Techniques such as heat maps, $t$-SNE, and decision trees help to visualise and explain how different modalities interact and contribute to outcomes, fostering user trust and supporting error analysis and model refinement.
Real-world deployment of multimodal AI systems
Deploying multimodal AI systems (the bottom of Fig. 2a) requires infrastructure for hosting, scaling, and ongoing management, with continuous monitoring to ensure the system remains effective and relevant paleyes2022challenges. Unlike unimodal AI, these systems face unique challenges, such as integrating diverse data modalities in real-time environments and maintaining performance in the face of modality-specific issues, such as sensor failures or data stream interruptions. Robust monitoring mechanisms enable the system to detect and compensate for failures in one modality by leveraging information from others. Additionally, real-time multimodal data fusion, where multimodal data may arrive asynchronously, requires careful infrastructure design and adaptive algorithms.
The following sections examine three cross-disciplinary use cases that demonstrate the value and applicability of the deployment-centric multimodal AI workflow: pandemic response, self-driving car design, and climate change adaptation.
Use case 1: Pandemic response
Figure 3: Deployment-centric multimodal AI workflow for the pandemic response use case. This example illustrates how the proposed deployment-centric multimodal AI workflow can better address real-world challenges in pandemic response, drawing on multidisciplinary collaboration, as indicated by the different colours across stages. The process begins with defining the problem of interest properly and carefully considering the space of various deployment constraints in order to formulate specific tasks. Two tasks are shown in the figure: one with a focus on healthcare and the other on social science. Next, we develop multimodal AI systems using diverse data sources supported by multidisciplinary collaboration. Finally, we proceed to real-world deployment and assessment, exemplified by the application of patient surge management.
Pandemic response (Fig. 3) presents a complex challenge requiring coordinated efforts across healthcare and other disciplines khojasteh2022climate. Defining the problem involves assessing early how multimodal AI can offer advantages over unimodal approaches by integrating diverse data sources for deeper insights and more accurate predictions. For example, in social science, combining textual content from surveys or social networks with tabular demographic data can help understand mental health impacts and detect changes in online behaviour elmer2020students; valdez2020social. In sustainability, analysing time-series environmental data, satellite imagery, and sustainability reports can reveal the environmental effects of pandemics liu2020near; zheng2020satellite. In finance, integrating economic indicators, transaction patterns, and market response data with health data can predict socio-economic impacts and inform public health strategies mosser2020central; nicola2020socio. A well-defined problem ensures clear constraints and tasks can be established.
The deployment of multimodal AI in pandemic response faces privacy, technical, and economic constraints. Privacy concerns arise from integrating diverse data sources such as EHRs, social media, and genomic information, which increases the risk of re-identification and amplifies the potential impact of data breaches. Maintaining public trust requires robust privacy-preserving techniques al2019privacy; Wendland:2022vo; che2023MultimodalFL and compliance with regulatory frameworks. Technical challenges include real-time monitoring, where multimodal AI systems must integrate and process data streams rapidly for timely interventions. The heterogeneity of health and social data complicates alignment and standardisation, particularly in international and/or multi-centre collaborations. Achieving minimum predictive accuracy thresholds for outbreak detection and public health decision-making helps build trust in AI-driven systems. Economic constraints, including computational costs and budget limitations, restrict accessibility and scalability. High-cost modalities (e.g. MRI) may be impractical in low-resource settings, so multimodal AI systems should flexibly support lower-cost alternatives (e.g. wearables) to promote equitable deployment.
Task formulation translates the defined problem into specific AI objectives, detailing inputs, outputs, and evaluation metrics for pandemic-related challenges. A primary task is predicting disease outbreaks by integrating multimodal data such as epidemiological records, EHRs, and environmental factors to inform early interventions and resource allocation. Another task is monitoring public adherence to health guidelines using social media data, geospatial information, and biosignals, providing actionable insights to refine policies. Additionally, multimodal AI can support real-time decision-making in healthcare by combining patient data streams, including wearable biosignals, medical images, and clinical notes, for managing patient surges effectively ding2020wearable; charlton_wearable_2022. Task formulation ensures that multimodal AI systems align with technical and practical needs, strengthening pandemic response through actionable insights and tailored interventions.
The development of multimodal AI systems for pandemic response adapts a standard workflow of the five key steps outlined above. Data collection assembles diverse sources, such as EHRs, genomic data, and biosignals, to create comprehensive datasets for outbreak prediction and patient management. During curation, standardisation and alignment address challenges such as integrating time-sensitive data streams from biosignals and geospatial sources. Multimodal learning can employ self-supervised zong2024self and cross-domain yang2014cross techniques to capture cross-modal relationships effectively and integrate domain-specific knowledge. Evaluation requires multi-metric, multi-centre validation han2024randomised; plana2022randomized, ensuring robustness for tasks such as outbreak prediction and public adherence monitoring. Interpretability and explainability enable healthcare professionals and policymakers to trust and act on multimodal AI insights in high-stakes scenarios imrie2023multiple. Tailoring these development steps to pandemic-specific needs ensures adaptability, reliability, and ethical compliance in rapidly changing contexts.
Deploying multimodal AI systems in pandemic response translates research into actionable solutions for high-stakes environments. For instance, patient surge management can integrate biosignals, medical images, and clinical notes to optimise intensive care unit bed allocation and staff deployment. Public health decision-making can benefit from multimodal analyses of epidemiological data, EHRs, and social media trends to guide interventions such as lockdowns or resource distribution. Real-time deployment requires robust infrastructure to process asynchronous data streams and address modality-specific interruptions, such as gaps in biosignal or social media data. Adaptive algorithms and reliable systems ensure timely and accurate outbreak prediction and monitoring. Assessment through multi-centre trials and multi-disciplinary benchmarks mincu2022developing can validate the added value of multimodal AI over unimodal systems. Measuring the accuracy of outbreak prediction or the effectiveness of intervention informed by multimodal AI can ensure trust and accountability. Further addressing scalability and privacy compliance can support sustainable and ethical deployment.
Use case 2: Self-driving car design
Self-driving car design (Supplementary Fig. S3) exemplifies a transformative application of multimodal AI, addressing challenges in safety, efficiency, and sustainability soni2021design; badue2021self. Defining the problem involves determining how multimodal AI can significantly outperform unimodal approaches for navigating complex environments safely and efficiently. For instance, fusing sensors such as LiDAR (light detection and ranging), radar and cameras can enhance perception and decision-making in dynamic settings, especially where a single sensor may fail yeong2021sensor. Beyond perception, multimodal AI can support physical simulations, such as combining text, images, and videos, to model vehicle dynamics and road interactions yang2024generalized. In materials design, integrating structural and chemical data aids in developing lightweight, sustainable materials for vehicle construction zhang2022advanced. Clearly defining these challenges ensures multimodal AI solutions are effectively tailored to the demands of autonomous driving.
The deployment of multimodal AI in self-driving car design faces safety, privacy, technical and economic constraints that must be addressed for safe and scalable adoption. Privacy concerns arise from integrating multimodal data, such as LiDAR, cameras and passenger-related information, which increases the potential for data misuse or unauthorised tracking hansson2021self. Mitigating these risks requires robust data anonymisation and adherence to privacy regulations chowdhury2020attacks. Technical challenges include real-time data processing in dynamic environments, such as urban intersections or adverse weather. Modality-specific failures, such as LiDAR disruptions in heavy rain, require fallback mechanisms such as radar to maintain reliability. Self-driving systems must achieve high predictive accuracy for collision avoidance and route planning, and operate reliably across regions with varying regulations and infrastructure. Economic constraints involve balancing the cost of high-resolution sensors with scalability needs. Meeting sustainability goals requires energy-efficient operation, with trade-offs between sensor performance and power consumption. Addressing these constraints holistically enables robust and accessible deployment of self-driving technologies.
In self-driving car design, task formulation maps problems and constraints to concrete AI objectives for perception, navigation, and safety. A key task is developing AI driving systems that integrate multimodal data streams, including LiDAR, radar, GPS, and cameras, for real-time navigation. These systems must dynamically process inputs to handle traffic, obstacles, and weather variability. For example, integrating road and weather data can enhance route planning and safety in adverse conditions dey2014potential. Another task is predicting material properties and performance to support vehicle design. By fusing structural, experimental, and supply chain data, multimodal AI can optimise materials for lightweight, durable, and cost-effective vehicles, aligning with sustainability goals kamran2022artificial. Additional tasks include passenger-focused objectives, such as integrating biosignals and cabin sensors to enhance comfort and safety. Ensuring privacy compliance remains integral across all tasks.
The development of multimodal AI systems for self-driving cars follows the five-step workflow outlined above, adapted to the demands of autonomous systems. Data collection involves assembling sensor inputs (e.g. LiDAR, radar, cameras) and materials data to build comprehensive datasets. Data curation addresses challenges such as synchronising asynchronous sensor inputs and standardising materials information. Learning can employ fusion strategies to enhance system reliability, while physics-informed neural networks raissi2019physics; karniadakis2021physics can integrate scientific laws to improve realistic simulations of driving scenarios and materials development. Evaluation ensures robustness through multi-metric validation across diverse scenarios, including urban environments and extreme weather conditions. Finally, interpretation in autonomous systems supports real-time explainability and transparency, enabling stakeholders to trust and audit system decisions. By aligning these steps with the unique requirements of autonomous systems, multimodal AI can ensure adaptability, safety, and performance.
Real-world deployment of multimodal AI systems for self-driving cars requires live sensor integration, infrastructure readiness, and operational resilience across diverse environments. Robust in-vehicle computing and inter-vehicle communication enable reliable and coordinated performance at scale. Assessment validates multimodal AI’s added value through simulations and multi-centre trials. Testing performance under extreme conditions, such as adverse weather or high-traffic intersections, ensures reliability. Addressing scalability and privacy compliance further supports ethical, sustainable deployment, advancing societal trust in self-driving technologies liu2019public.
Use case 3: Climate change adaptation
Climate change adaptation (Supplementary Fig. S3) involves forecasting extreme weather, assessing climate risks, and managing resources to minimise environmental and socio-economic impacts bi_accurate_2023; lam2023learning; mathiesen2018rating; breitenstein2022disclosure. Defining the problem involves assessing whether multimodal AI can surpass unimodal approaches by integrating diverse data sources for deeper insights and actionable strategies. For example, multimodal AI can combine satellite imagery, historical climate records, and geospatial data to model weather patterns and predict extreme events, supporting proactive disaster management imran2020using. Integrating socio-economic data with environmental observations can enable more comprehensive climate risk assessments tanir2024social.
Deploying multimodal AI for climate change adaptation must navigate privacy, technical and economic constraints. Privacy concerns arise from integrating sensitive data, such as socio-economic records and proprietary satellite imagery. Ensuring compliance with global and regional privacy regulations and the ethical use of data supports responsible deployment. Technical challenges include the heterogeneity and sparsity of environmental and socio-economic data. Economic constraints, such as the high cost of advanced sensors, necessitate balancing low-cost and low-fidelity data sources with high-cost, high-fidelity alternatives to maintain accessibility and scalability. By addressing these constraints, multimodal AI can deliver equitable and sustainable climate solutions in diverse global contexts.
For climate change adaptation, task formulation aligns complex environmental challenges with actionable AI tasks. A primary task is forecasting extreme weather events by integrating multimodal data such as satellite imagery, historical climate records, and real-time observations. These forecasts can enable timely interventions to mitigate human and economic losses kumar2021overview, including early warning systems for floods that integrate geospatial, hydrological, and social data streams tkachenko2017predicting. Another task is predicting the socio-economic impacts of severe weather by combining spatial, market, and socio-economic data, guiding resource allocation and policy decisions thulke2024climategpt. These tasks align multimodal AI with both technical requirements and societal priorities, advancing climate resilience and equitable adaptation strategies.
Developing multimodal AI systems for climate change adaptation follows a similar five-step workflow. Data collection gathers remote sensing data, near-real-time sensor network outputs, historical climate records, and socio-economic metrics, creating comprehensive datasets for tasks such as extreme weather forecasting and early warning systems. Data curation addresses challenges such as temporal misalignment and data sparsity by standardising inputs and enriching sparse data through interpolation or integration of auxiliary sources. Multimodal learning can benefit from advanced models such as Aurora bodnar2024aurora, which fuse satellite imagery and meteorological data to improve atmospheric predictions. Evaluation uses multi-metric benchmarks to test system performance in diverse scenarios. For example, WeatherBench 2 provides a standard platform for assessing atmospheric prediction models rasp2024weatherbench. Interpretation ensures outputs are transparent and actionable for stakeholders, supporting evidence-based decision-making. By aligning these steps with climate-specific challenges, multimodal AI can deliver reliable, scalable, and adaptable solutions.
Real-world deployment of multimodal AI for climate change adaptation translates research insights into scalable, operational systems. Early warning platforms and impact forecasting models must synchronise diverse data streams, deliver real-time outputs, and integrate with decision-making infrastructures across sectors and regions morshed2024decoding; Reichstein2025EarlyWO. Real-time monitoring requires robust infrastructure to synchronise and process diverse modalities. Deployment challenges include latency, data coverage gaps, and variable infrastructure capacity across geographies. Assessment involves stress testing and multi-centre validation to ensure reliability across environmental and socio-economic scenarios. Ethical considerations, such as prioritising vulnerable communities and ensuring equitable access to data and AI tools, help enable sustainable deployment. By overcoming deployment challenges, multimodal AI systems can empower policymakers, industries, and communities to respond proactively to climate challenges, advancing global sustainability efforts.
Outlook
Most existing research on multimodal AI has focused on model-centric development. While attention to data-centric development is increasing, unlocking the full potential of multimodal AI and addressing real-world challenges will ultimately require a shift towards deployment-centric development. This shift will require strategic advancements in data, model, and deployment methods, as well as concerted efforts to address multimodal-AI-specific challenges, namely modality incompleteness, multimodal heterogeneity, cross-modality alignment, modality complementarity, and multimodal privacy risk, through interdisciplinary collaborations and community-building. Box 1 presents strategic recommendations to advance multimodal AI across disciplines under a deployment-centric perspective. These recommendations draw on insights from the three real-world use cases discussed above and the broader cross-disciplinary analysis in the Supplementary Information on multimodal AI beyond the three use cases, providing actionable guidance for future developments.
Deployment-centric development brings challenges around safety, reliability, interpretability, scalability, and ethics, particularly as multimodal AI expands in high-stake applications like healthcare and sustainability. Addressing these challenges requires implementing robust human-in-the-loop systems, developing clear standards for safety and transparency, and prioritising scalable, resource-efficient infrastructure to support increasing demands for data integration, processing, and storage.
Robust data-centric development underpins deployment-centric progress by ensuring data availability, diversity, and quality. Quality multimodal data is often limited, and existing datasets often lack the diversity needed to ensure fairness in AI systems. Clear benchmarks and globally accessible datasets are needed to enhance trust, consistency, and adaptability across disciplines, facilitating reproducible research (Supplementary Fig. S1). Promoting secure data-sharing frameworks torabi2023common, open initiatives like the European Life Science Infrastructure for Biological Information crosswell2012elixir, and globally accessible multimodal data platforms will further strengthen AI’s capacity to address diverse global challenges. Knowledge graphs liang2024survey also offer promise for organising varied data formats to unify and contextualise multimodal inputs across disciplines.
Model-centric development for multimodal AI faces unique challenges in effectively and efficiently fusing diverse modalities. While foundation models have proven effective for vision and language tasks, expanding them to other modalities and disciplines archit2025segment could significantly lower development barriers. Guidelines and frameworks for selecting relevant modalities, comparing multimodal versus unimodal performance, and evaluating the benefits of using many versus few modalities will help maximise model efficiency and utility.
Interdisciplinary collaboration is crucial to progress, as discussed in the context of the three use cases. Standardising data integration practices, fostering cross-disciplinary knowledge exchanges, and establishing shared platforms will help bridge disciplinary divides, enabling multimodal AI to address complex societal challenges holistically.
Building a dynamic and inclusive multimodal AI community fosters innovation and drives solutions to complex real-world challenges. Collaborative efforts, through regular workshops, forums, and interdisciplinary research initiatives, facilitate knowledge exchange and inspire collective problem-solving across disciplines. Engaging researchers, practitioners, and stakeholders from diverse backgrounds, particularly early-career researchers and underrepresented groups, will not only enrich the AI ecosystem but also broaden the perspectives and expertise shaping it, making it more ‘multimodal’. Community-driven initiatives Chan2020ReportingGF; de2022guidelines, including comprehensive surveys or perspective papers, provide insights that guide future developments, ensuring multimodal AI advances in responsible, impactful directions.
Ultimately, the success of multimodal AI lies in its deployment. By embracing a deployment-centric mindset, the community can turn research breakthroughs into real-world impact across disciplines.
Box 1    Strategic recommendations for advancing multimodal AI across disciplines Deployment-centric development Safety, reliability, and interpretability Challenge: Ensuring reliable, safe deployment across varied conditions and building user trust to facilitate understanding and broader adoption. Recommendation: Translate complex data and outputs into accessible formats (images, text, or speech); incorporate domain-specific knowledge; implement human-in-the-loop systems for verification and validation; conduct rigorous usability testing; develop standards for safety, reliability, and interpretability criteria; and enforce such criteria in research and peer review. Scalability and resource efficiency Challenge: Scaling multimodal AI to handle vast data volumes while optimising resource consumption, maintaining system availability, and ensuring uninterrupted communication between system components. Recommendation: Innovate scalable and resource-efficient multimodal AI solutions; promote collaboration between AI developers and hardware providers for sustainable resource use; ensure stable system availability and seamless inter-component communication; and develop cloud-based solutions and edge computing to power scalability. Ethical compliance and user preparedness Challenge: Addressing privacy, consent, and bias in applications with significant societal impacts. Recommendation: Develop clear ethical guidelines; include end-users in the design; involve human oversight and create user feedback loops for continuous solution refinement; and provide comprehensive training for reliable deployment. Data-centric development Data scarcity and access Challenge: Limited availability of high-quality multimodal datasets spanning a comprehensive range of modalities, complicated further by privacy and ethical constraints and additional heterogeneity when data sharing spans secure data environments, organisations, and regions. Recommendation: Promote open data initiatives, such as anonymised data-sharing programmes; invest in cross-discipline efforts for data collection and curation to improve availability; build data infrastructures that ensure secure, privacy-compliant, and ethics-compliant access across domains and regions; establish guidelines to standardise formats and annotations; and develop advanced tools such as knowledge graphs to integrate structured and unstructured data sources for improved contextual understanding, reasoning, and interoperability. Balanced data representation Challenge: Bias due to the limited availability of data representing underrepresented or less-studied categories (e.g. populations and regions). Recommendation: Diversify dataset development to enhance model generalisability and reduce bias; and build and enhance multimodal data platforms (e.g. UK Biobank bycroft2018uk, MIMIC johnson2016mimic, Materials Project jain2013commentary, ERA5 hersbach2020era5) to improve global data accessibility and quality. Model-centric development Multimodal fusion and modality selection Challenge: Managing diverse properties and patterns across modalities, including avoiding unnecessary modality redundancy, to maintain integrity and utility while enhancing integration. Recommendation: Assess the value of additional modalities by comparing unimodal, few-modality, and full-modality models; enhance semantic modality alignment, e.g. via large language models; develop guidelines for optimal modality selection; and address other multimodal-AI-specific modelling challenges (Fig. 2b). Foundation models (FMs) Challenge: Developing FMs beyond vision and language requires substantial resources, limiting accessibility across disciplines. Recommendation: Develop and expand multimodal FMs beyond vision and language as critical reusable infrastructure to lower barriers; and invest in downstream research and advance methods to reduce resource costs. Interdisciplinary collaboration Cross-disciplinary standards and communication Challenge: Discrepancies in multimodal data standards and communication barriers across disciplines. Recommendation: Promote cross-disciplinary standards for multimodal data integration; and organise interdisciplinary exchange events to align goals, foster collaboration, and build a thriving ecosystem. Intellectual property (IP) and workflow adaptation Challenge: IP concerns and resistance to change established workflows. Recommendation: Develop adaptive strategies and open collaborative platforms to facilitate cross-disciplinary projects.
Data availability
Source data for Fig. 1 and Fig. S2 is available with this paper and updated annually at https://github.com/multimodalAI/multimodal-ai-landscape.
Acknowledgements
[…]
Author contributions
X.L., J.Z., S.Z. and H.L. contributed equally. X.L., J.Z., S.Z., T.L.vdP., S.T., T.Z., W.K.C., P.H.C., and H.L. conceptualised the manuscript. X.L., J.Z., S.Z., A.G., M.Z., L.vZ., O.T., and H.L. designed the figures and table. X.L. and H.L. coordinated the entire project. All authors contributed to writing, resources, or editing.
Competing interests
G.G.S. is a scientific advisory board member at BioAIHealth. P.H.C. provides consulting services to Cambridge University Technical Services for wearable manufacturers. The remaining authors declare no competing interests.
Supplementary Information
Towards deployment-centric multimodal AI beyond vision and language
S1. Multimodal AI beyond the three use cases
This section highlights discipline-specific advancements and challenges in multimodal AI that extend beyond the three primary use cases of pandemic response, self-driving car design, and climate change adaptation. By exploring advancements and challenges in healthcare, social science, engineering, science, sustainability, and finance, we showcase the versatility of multimodal AI across diverse domains. These insights underline the importance of interdisciplinary collaboration in addressing complex global challenges while paving the way for future advancements.
Healthcare
Beyond its critical role in pandemic response, multimodal AI holds transformative potential in healthcare by advancing diagnostics, personalising treatment, and improving patient care. By integrating diverse data sources such as electronic health records (EHRs), imaging, biosignals, and multiomics, these systems can provide comprehensive insights into patient health acosta2022multimodal. For example, integrating wearable biosignals with EHRs can enable continuous monitoring, allowing for the early detection of chronic conditions such as arrhythmias and heart failure krittanawong2021integration. In oncology, the fusion of imaging modalities and multiomics data can enhance cancer diagnosis and treatment planning steyaert2023multimodal, offering a more nuanced understanding of disease progression. Similarly, in neurology, combining imaging, sensor data, and clinical observations can facilitate the early detection of neurodegenerative diseases huang2023multimodal.
Wider adoption of multimodal AI is hindered by challenges such as privacy concerns, the heterogeneity of healthcare data, and the scarcity of high-quality datasets. Addressing these challenges requires standardised methodologies, privacy-preserving techniques, and interdisciplinary collaborations. Building datasets that reflect the complexities of patient pathways and addressing bottlenecks in data integration are crucial steps towards integrating multimodal AI into routine clinical practice. Techniques like federated learning, differential privacy, and privacy-preserving generative AI for synthetic multimodal data generation can mitigate privacy risks, enabling the development of robust and trustworthy AI systems. Integrating multimodal data, such as multiomics lunke2023integrated, into routine clinical workflows will require rigorous validation and improved standardisation. Notably, limited validation through randomised controlled trials han2024randomised has led to the emergence of community-based guidelines Chan2020ReportingGF aimed at improving AI reliability and transparency.
Social science
In social science, multimodal AI can drive innovations in behavioural analysis, public policy decision-making, and societal impact assessments. Beyond its applications in pandemic response, multimodal AI can help monitor societal behaviours and safeguard vulnerable populations, such as protecting children online, by integrating text, image and metadata from social media ali2023getting. Multimodal AI systems can also help detect online criminal activity goyal2023detection and analyse societal trends to inform public policy androutsopoulou2018framework. Additionally, acoustic data can add depth to communication analysis by capturing emotional nuances zhu2024review, while combining economic transaction data with geospatial information can reveal trends in societal decision-making Huang2024CrowdsourcedGD.
Persistent challenges include restricted access to sensitive data such as social media or mental health records, privacy concerns, and the complexity of annotating subjective human behaviours that differ across cultures. Moreover, specialised evaluation metrics are needed to assess the effectiveness of these systems in capturing nuanced social trends. To advance multimodal AI in social science, efforts should focus on expanding data diversity, addressing privacy challenges, developing robust annotation and evaluation methods, and fostering robust ethical frameworks and cross-disciplinary expertise. These efforts will ensure that the insights generated are fair, transparent, and aligned with societal goals.
Engineering
Beyond self-driving cars, multimodal AI supports advanced autonomous systems such as robotics, precision manufacturing, and medical applications. Robots equipped with multimodal AI can integrate vision, haptic feedback, and sensor data to perform delicate tasks, such as in surgical procedures he2022robotic or complex manipulation wang2024multimodal, where precision is paramount. Additionally, multimodal AI can enable seamless human-robot interaction by combining speech recognition, visual data, and environmental context duncan2024survey. Speech technology and natural language processing can enhance user interfaces, improving interaction while minimising distractions for operators in some robotic applications.
Despite the potential of multimodal AI, challenges remain in achieving effective data fusion, system interoperability, and robust model generalisation. Furthermore, the lack of diverse datasets, particularly for edge cases or atypical deployment environments, limits the scalability of these systems. Establishing open platforms for data and toolkit sharing accelerates innovation and ensures interoperability across engineering applications. Integrating insights from legal studies and social sciences can enhance compliance with evolving regulations and improve user-centred design, boosting societal acceptance of autonomous systems. Environmental science can contribute to developing energy-efficient and sustainable designs, particularly for robotics and manufacturing processes. Addressing these challenges requires interdisciplinary collaborations to improve data diversity and develop scalable, interoperable systems.
Science
Multimodal AI is transforming scientific discovery by integrating diverse data sources to model complex phenomena. Beyond its role in vehicle dynamics simulation, as seen in self-driving car design, multimodal AI is accelerating materials science advancements. For instance, integrating experimental and computational data enables the development of improved batteries, fuel cells and supercapacitors liu2024recent. We can incorporate domain knowledge via techniques such as physics-informed neural networks raissi2019physics; karniadakis2021physics and neurosymbolic AI garcez2023neurosymbolic to improve the accuracy of scientific simulations and enhance model trustworthiness and predictive power across applications yang2023context; rezaei2021learning.
Remaining challenges of multimodal AI for scientific discovery lie in integrating heterogeneous data, ranging from atomic-level properties to macroscopic observations, and incorporating synthetic data while addressing the scarcity of high-fidelity datasets amidst an abundance of low-fidelity data. Advancing scientific discovery with multimodal AI requires comprehensive datasets, rigorous data standards, and interdisciplinary research environments. These efforts will ensure that models are aligned with real-world dynamics, unlocking new possibilities in materials science, physics, and beyond trinh2024olympiad; raayoni2021ramanujan.
Sustainability
Multimodal AI can help address sustainability challenges, such as biodiversity conservation and environmental monitoring. Beyond its applications in climate change adaptation, multimodal AI can combine geospatial data, bioacoustic recordings, and environmental DNA to track species and ecosystems pollock2025harnessing, supporting conservation planning and enhancing understanding of ecosystem dynamics. In terrestrial environments, ecoacoustic and LiDAR data can be combined to model biodiversity variation across complex landscapes rappaport2020acoustic, while in marine settings, integrating remote sensing with in situ sensors enables long-term monitoring of ecosystem health and debris pathways maximenko2021integrated.
Geographic imbalances in data availability, particularly in underrepresented regions, limit model fairness and accuracy. Marine environments pose unique sensing challenges briciu2023sensors, including high turbidity and limited optical visibility, which constrain conventional imaging approaches. This necessitates innovative solutions such as autonomous underwater vehicles and acoustic sensors to access deeper waters, remote habitats, and ecologically sensitive regions whitt2020future. Future advancements will depend on building diverse, scalable datasets, de-biasing models, and leveraging digital twins Li2023 for real-time monitoring and actionable insights. By addressing these challenges, multimodal AI can contribute significantly to global sustainability efforts zhao2024artificial.
Finance
Multimodal AI can advance finance by integrating financial, environmental, and social data to improve risk assessments, market forecasting, and sustainable investment strategies. In particular, it can enhance ESG (environmental, social, and governance) investing by helping stakeholders make more responsible financial decisions ang2024learning. Integrating financial data and environmental data such as satellite imagery can enable a more comprehensive approach to risk management and resilience leng2024power. Additionally, graph-based multimodal models can monitor systemic risks by capturing complex relationships within financial ecosystems ang2024learning, supporting the early detection of market disruptions. Traditional financial models often struggle to respond to economic shocks such as natural disasters. Integrating real-time, multimodal data can improve model adaptability and accuracy under volatile conditions, ultimately enhancing financial stability cao2023ai.
Challenges such as data privacy, regulatory complexity, and the need for explainability remain significant barriers. Addressing these issues requires comprehensive multimodal datasets and transparent evaluation frameworks to build trust and ensure compliance with regulations. Interdisciplinary collaborations can help refine multimodal AI systems in finance. Environmental science informs environmental risk assessments, enabling more resilient investment strategies and credit evaluations in the face of severe adverse events. Social sciences provide socio-political insights that promote transparency, equity, and ethical compliance, ensuring financial practices align with societal goals. These collaborations align financial systems with ethical and sustainable practices while equipping stakeholders to make resilient decisions in an evolving financial landscape.
Multimodal benchmark datasets
Figure S1: Examples of multimodal benchmark datasets. We selected two illustrative examples for each of the six disciplines to showcase diverse multimodal data and provide a starting point for multimodal AI exploration. For each dataset, we report six key attributes: year of release, data origin, data modalities, accessibility, scale of the primary modalities, and example applications. In each discipline, datasets are listed in ascending order by year of release. For the healthcare, science, and finance disciplines, we selected one small-scale dataset for quick experimentation and one larger-scale dataset for comprehensive exploration. The datasets include: TCGA (The Cancer Genome Atlas) tomczak2015review; MIMIC-IV (The Medical Information Mart for Intensive Care IV database) johnson2023mimic; Trafficking-10k (human trafficking advertisement dataset) tong2017combating; IGDD project (Instagram data donation project) razi2022instagram; nuScenes (autonomous driving scene dataset) caesar2020nuscenes; DAIR-V2X (real-scenarios vehicle to everything dataset) yu2022dair; Materials project (material property dataset) jain2013commentary; MathVista (mathematical reasoning dataset) lu2024mathvista; iNaturalist (citizen science platform for biodiversity data) inaturalist; ERA5 (European Centre for Medium-Range Weather Forecasts Reanalysis v5) hersbach2020era5; Monetary policy calls (monetary policy call dataset) mathur2022monopoly; China A-shares market (dataset of public companies listed in China A-shares market) cheng2022financial. IGDD and ERA5 are not multimodal in the traditional sense but are treated as such under our broader definition.
Benchmark datasets are crucial for developing multimodal AI systems. Figure S1 illustrates examples of selected benchmark datasets across the six discussed disciplines, serving as valuable resources for experimental exploration and evaluation. Each dataset is described by its year of release, data origin, data modalities, accessibility, scale of the primary modalities, and example applications. For the healthcare, science, and finance disciplines, we selected one small-scale dataset for quick experimentation and one larger-scale dataset for comprehensive exploration. In our selection, we prioritise datasets having multiple modalities, particularly those beyond vision and language, and with free and open access, although this is not always possible. For example, the Instagram data donation (IGDD) project and Trafficking-10k datasets require special permissions or approvals. In addition, there are multiple ways to define and categorise the modalities within a dataset, so the modalities listed in Fig. S1 represent just one possible approach. For example, the IGDD dataset contains only text data, and we treat its two subtypes, chat histories and surveys, as distinct text modalities, following our broader definition of multimodality. The ERA5 dataset comprises spatiotemporal data, and we consider variables such as temperature, relative humidity, and wind velocity as multiple ‘views’ (in the broader sense defined in this Perspective), each offering distinct information for weather forecasting and climate research, as in previous studiesjin2023spatiotemporal; chen2023fengwu.
S2. Methods and extended analysis of the multimodal AI landscape
Figure S2: Landscape of multimodal AI research from 2019 to 2024. a, Yearly growth in the number of multimodal AI preprints on arXiv (bars) and their proportion among all AI preprints (line), demonstrating a marked rise from 2019 to 2024. b, Distribution of multimodal AI preprints by the number of combined modalities. The inset line plot shows the growth rate of each combination. c, Detailed breakdown of pairwise, triple, and quadruple modality combinations, demonstrating the dominance of pairwise combinations and the increasing prevalence of more complex combinations. d, Heatmap of combinations of ‘other’ modalities (excluding vision and language) by year. Each row in the heatmap represents the number of preprints published in a given year, with darker shades indicating higher counts. The coloured circles on the left correspond to the modality combinations for each row. For example, in 2023, there were only two preprints using graph and tabular data (the cell in the sixth row and the fifth column).
All arXiv preprints ginsparg2011arxiv from 2019 to 2024 were analysed to examine the latest trends in multimodal AI research. ArXiv was chosen for its open access, broad adoption within the AI community, and ability to reflect emerging developments more rapidly than peer-reviewed platforms. Raw data were extracted from the publicly available arXiv metadata dump arxiv_org_submitters_2024 hosted on Kaggle and filtered using queries for six general AI terms “AI”, “A.I.”, “artificial intelligence”, “machine learning”, “deep learning”, and “neural network” appearing in either the title or the abstract to identify relevant preprints. The list was further refined to identify multimodal AI preprints by searching for “multimodal” and “multi-modal”, and specific modalities were identified through targeted queries: “vision”, “image”, “video”, and “visual” for vision; “text”, “language”, and “textual” for language; and similar sets of terms for the other six modalities: time series, graph, audio, spatial, sensor, and tabular data (see Data availability). As these queries are inherently approximate, conclusions are confined to overall trends supported by strong evidence rather than specific numerical details.
Figures 1 and S2 summarise the findings. Figure 1 shows the dominance of vision and language in multimodal AI research. Figure S2a depicts an accelerating upward trend in both the number and proportion of multimodal AI preprints since 2022, highlighting the field’s rapid expansion. Figure S2b categorises multimodal AI preprints by the number of modalities and their respective growth rates for additional insights. Pairwise combinations are the most popular, while combinations involving more modalities are steadily gaining attention, supporting richer data fusion for tasks requiring diverse information sources. Figure S2c examines pairwise, triple, and quadruple modality combinations across four modality clusters: vision and language, vision and others, language and others, and others. The combination of vision and language remains dominant, particularly in triple and quadruple combinations. In particular, from 2023 to 2024, pairwise combinations involving language have grown significantly, driven by the revolution in large language models. Finally, Fig. S2d details underexplored modality combinations beyond vision and language, ordered by their 2024 preprint counts. Time series and sensor are the most popular non-vision, non-language modality combinations, likely due in part to the fact that sensor data are often recorded as time series. The popularity of the time series and spatial combination may be partially attributed to the importance of spatiotemporal modelling. In contrast, non-vision, non-language modality combinations involving graph, audio, and tabular data are much less explored, highlighting their untapped potential in multimodal AI research.
Figure S3: Deployment-centric multimodal AI workflow for the self-driving car design and climate change adaptation use cases. These examples are depicted in a similar way as Fig. 3, showing the diverse data sources needed and the multidisciplinary collaboration involved.
S3. Illustration of the self-driving car design and climate change adaptation use cases
Figure S3 illustrates the deployment-centric workflow for the self-driving car design and climate change adaptation use cases. In self-driving car design (top), the workflow incorporates deployment constraints from social science, engineering, sustainability, and finance disciplines. Two key tasks focus on engineering and science considerations, respectively. Multimodal AI development uses multiple data sources, including remote sensing, weather, and materials data, aiming for real-world deployment exemplified by automated intersection turning. In climate change adaptation (bottom), the workflow considers constraints such as mass deployment cost barriers, minimum data sampling density, and prioritisation of vulnerable regions. Task formulation and multimodal AI development focus on environmental sustainability and financial considerations, reflecting how task and model development are shaped by discipline-specific needs. Early flood warning serves as an example of real-world deployment and assessment. Together, Fig. 3 and Fig. S3 demonstrate the versatility of the proposed workflow, showing how deployment-centric considerations can be applied across different disciplines to solve complex challenges.